# RobustSplat: Decoupling Densification and Dynamics for Transient-Free 3DGS

### v2

@@ 3.1 Overview
To better leverage deep high-dimensional feature information extracted from images, which have different properties as the image residual, we adopt a feature robust loss $\mathcal{L}_{\textrm{cos}}$ utilizing the information of feature similarity between the rendered and captured images. Specifically, we extract DINOv2 features of the real image $f_{t}$ and rendered image $f_{t}^{\prime}$, and compute their cosine similarity map. Then we convert the cosine similarity map to be in the value range of $[0,1]$ following [18]:
@@
where $M_{\textrm{cos}}$ will be $1$ is the feature cosine similarity is 1, and it will be 00 if the similarity is less than $0.5$. Then the feature robust loss is expressed as:
@@ 4 Experiments
We evaluate our RobustSplat on two challenging benchmark datasets: NeRF On-the-go [31] and RobustNeRF [33]. The NeRF On-the-go dataset consists of a total of 12 scenes, featuring varying occlusion levels ($5\%\sim 30\%$). Among them, six scenes are widely used, while the remaining six are more complex scenarios, referred to as NeRF On-the-go II in this paper.
@@
In the mask bootstrapping, the lowest spatial resolution features are extracted from images of size $(224\times 224)$, while the highest spatial resolution features are derived from size $(504\times 504)$. Following existing methods [34], we apply a downsampling factor of 8 on the NeRF On-the-go and RobustNeRF datasets (factor 4 for specific scenarios, e.g., arcdetriomphe and patio). Low-resolution residuals are further downsampled by an additional factor of 4 based on this configuration.
@@
We evaluated our RobustSplat against multiple baselines, including the vanilla 3D Gaussian Splatting [16] which we built upon, and recent robust methods including SpotLessSplat [34], WildGaussians [18], Robust3DGS [39] and T-3DGS [30]. To ensure a fair comparison, we utilized the publicly available implementations of these methods and conducted evaluations using the same camera matrices across all experiments. We assessed performance through both visual comparisons and quantitative metrics, employing PSNR, SSIM, and LPIPS to measure reconstruction quality.
@@ Appendix B More Details for the Method
The original 3DGS  [16] resets the opacity starting from the 3000 iterations while maintaining an interval of 3000 iterations. This operation aims to eliminate the accumulation of low-opacity Gaussian primitives in regions close to the camera, which can interfere with gradient backpropagation and manifest as artifacts. However, the opacity reset is no longer suitable for our method due to the delayed Gaussian growth. Therefore, we delay the opacity reset to start from the 15000 iterations while maintaining the same interval of 3000 iterations.
@@
Appendix C Evaluation on On-the-go II Dataset
@@
Appendix D Comparison of Mask Estimation

### v3

@@ 3.1 Overview
To better leverage deep high-dimensional feature information extracted from images, which have different properties as the image residual, we adopt a feature robust loss $\mathcal{L}_{\textrm{cos}}$ utilizing the information of feature similarity between the rendered and captured images. Specifically, we extract DINOv2 features of the real image $f_{t}$ and rendered image $f_{t}^{\prime}$, and compute their cosine similarity map. Then we convert the cosine similarity map to be in the value range of [0,1][0,1][ 0 , 1 ] following [18]:
@@
where $M_{\textrm{cos}}$ will be 111 is the feature cosine similarity is 1, and it will be 0 if the similarity is less than 0.50.50.5. Then the feature robust loss is expressed as:
@@ 4 Experiments
[…]
@@
In the mask bootstrapping, the lowest spatial resolution features are extracted from images of size (224×224)(224$\times$ 224)( 224 × 224 ), while the highest spatial resolution features are derived from size (504×504)(504$\times$ 504)( 504 × 504 ). Following existing methods [34], we apply a downsampling factor of 8 on the NeRF On-the-go and RobustNeRF datasets (factor 4 for specific scenarios, e.g., arcdetriomphe and patio). Low-resolution residuals are further downsampled by an additional factor of 4 based on this configuration.
@@
We evaluated our RobustSplat against multiple baselines, including the vanilla 3D Gaussian Splatting [16] which we built upon, and recent robust methods including SpotLessSplats [34], WildGaussians [18], Robust3DGaussians [39] and T-3DGS [30]. To ensure a fair comparison, we utilized the publicly available implementations of these methods and conducted evaluations using the same camera matrices across all experiments. We assessed performance through both visual comparisons and quantitative metrics, employing PSNR, SSIM, and LPIPS to measure reconstruction quality.
@@ 5 Conclusion
Acknowledgements
This work was in part supported by the National Key R&D Program of China (Grant No. 2022ZD0119200), NSFC (Grant Nos. 62202409, 62472453), Guangdong Natural Science Foundation (No. 2025A1515010782), Shenzhen Science and Technology Program (No. JCYJ20220818102012025), CIE-Smartchip research fund (No. 2024-08), the Key Technology Project of Shenzhen (Grant No. KJZD20230923115104009), and Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001).
@@ Appendix B More Details for the Method
The original 3DGS  [16] resets the opacity starting from the 3000 iterations while maintaining an interval of 3000 iterations. This operation aims to eliminate the accumulation of low-opacity Gaussian primitives in regions close to the camera, which can interfere with gradient backpropagation and manifest as artifacts. However, the opacity reset is no longer suitable for our method due to the delayed Gaussian growth. Therefore, we delay the opacity reset to start from the 15000 iterations while maintaining the same interval of 3000 iterations. Meanwhile, the start of pruning is also delayed to 10000 iterations to align with delayed Gaussian growth.
@@
Appendix C Runtime Evaluation
Our method adopts the lightweight DINOv2 model ViT‑S/14‑distilled, with a feature dimensionality of 384, for feature extraction. As shown in Table S1, our method runs slightly slower than 3DGS but remains faster than other methods. SpotLessSplats achieves similar optimization time without iterative feature extraction, but its SD features, with a dimensionality of 1280, require a long processing time before training.
Table S1: Runtime evaluation on an NVIDIA RTX 3090 (unit: minutes). The runtime of SpotLessSplats is divided into two parts: training and SD feature extraction.
Method Mountain Fountain Corner Patio Spot Patio-High
#Img 120 #Img 169 #Img 101 #Img 99 #Img 169 #Img 222
3DGS 12.21 14.37 9.986 7.707 11.68 12.82
SpotLessSplats 13.48+6.9 16.07+9.8 14.15+6.4 13.82+6.4 13.03+9.5 14.07+13.7
WildGaussians 32.63 52.90 33.58 29.93 27.32 33.86
Ours 15.43 17.33 13.32 12.82 12.95 14.35
Appendix D More Ablation Study
Effects of Mask Regularization.
Initial mask estimation yields suboptimal results in most scenes due to unconverged reconstruction at early training stages. To address this challenge, we introduce a mask regularization for stabilizing early-stage mask training. Table S2 shows that removing the proposed mask regularization leads to a decrease in overall performance.
Table S2: Effects of Mask Regularization. We denote Mask Regularization as “MR”.
Method Mountain Fountain Corner Patio Spot Patio-high
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Ours w/o MR 21.09 0.728 20.87 0.701 26.18 0.889 21.61 0.826 25.63 0.907 22.68 0.837
Ours 21.15 0.737 21.01 0.701 26.42 0.897 21.63 0.827 26.21 0.907 22.87 0.837
Effects of Delayed Gaussian Growth.
We discussed the effectiveness of Delayed Gaussian Growth in Section 4.3 of the main paper. To further validate its effects, we extend the Delayed Gaussian Growth to 3DGS in this supplementary material. Table S3 shows that integrating the delayed Gaussian growth into 3DGS leads to improve results, but its performance is limited by the lack of predicting the transient masks.
Table S3: Effects of Extending Delayed Gaussian Growth to 3DGS. We denote Delayed Gaussian Growth as “DG”.
Method Mountain Fountain Corner Patio Spot Patio-high
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
3DGS 19.21 0.691 20.08 0.686 22.65 0.835 17.04 0.713 18.54 0.717 17.04 0.657
3DGS+DG 20.14 0.693 20.35 0.683 23.54 0.864 17.46 0.728 23.42 0.854 18.87 0.728
Ours 21.15 0.737 21.01 0.701 26.42 0.897 21.63 0.827 26.21 0.907 22.87 0.837
Appendix E Evaluation on On-the-go II Dataset
@@
Appendix F Comparison of Mask Estimation